\begin{thebibliography}{99}
		
		
		%		\bibitem{AB}
		%		Alazard  T., Baldi P.,
		%		\emph{Gravity capillary standing water waves},
		%		Arch. Rat. Mech. Anal. 217(3), 741-830, 2015.
		
%		\bibitem{ADM}
%		Antonelli, P.; Dolce, M.; Marcati, P.:
%		\emph{Linear stability analysis for 2D shear flows near Couette in the isentropic compressible Euler equations}.
%		Ann PDE 7, 24, (2021). 

		\bibitem{Baldi24}
		Baldi, P.:
		\emph{Nearly toroidal, periodic and quasi-periodic motions of fluid particles driven by the Gavrilov solutions of the Euler equations}.
		J. Math. Fluid Mech. 26:1, (2024).
		
		\bibitem{BBHM} 
		Baldi P.; Berti M.; Haus E.; Montalto R.: 
		\emph{Time quasi-periodic gravity water waves
			in finite depth}.
		Invent. Math. 214 (2), 739--911, (2018).
		
		%		\bibitem{BBM14}
		%		Baldi P., Berti M., Montalto R.,
		%		\emph{KAM for quasi-linear and fully nonlinear forced perturbations of Airy equation},
		%		Math. Ann. 359, 471–536, 2014.
		
		\bibitem{BM20}
		Baldi P.; Montalto R.:
		\emph{Quasi-periodic incompressible Euler flows in 3D}.
		Advances in Mathematics 384, 107730, (2021).
		%		
%		\bibitem{BaBM}
%		Bambusi D.; Berti M.; Magistrelli E.:
%		\newblock {\it Degenerate KAM theory for partial differential equations}. 
%		\newblock Journal Diff. Eq. 250(8), 3379--3397, (2011).
%		
%		\bibitem{BeMa}
%		Bedrossian J.;  Masmoudi N.: 
%		\emph{Inviscid damping and the asymptotic stability of planar shear
%			flows in the 2D Euler equations}. 
%		Publ. Math. Inst. Hautes Études Sci. 122, 195--300, (2015).
%		
%		\bibitem{BertiUMI}
%		Berti, M.: 
%		\emph{KAM for PDEs}.
%		Boll. Unione Mat. Ital. 9:115--142, (2016).
%		
%		\bibitem{BB}
%		Berti M.; Bolle P.:
%		\emph{A Nash-Moser approach to KAM theory}.
%		Fields Institute Communications, special volume "Hamiltonian PDEs and Applications", pp. 255--284, (2015).
		
		\bibitem{BFM}
		Berti M.; Franzoi L.; Maspero A.:
		\emph{Traveling quasi-periodic water waves with constant vorticity}. Arch. Rat. Mech. Anal. 240, 99--202, (2021). 
		
		\bibitem{BFM21}
		Berti M.; Franzoi L.; Maspero A.:
		\emph{Pure gravity traveling quasi-periodic water waves with constant vorticity}.  
	Comm. Pure Appl. Math. 77(2), 990--1064, (2024).
		%		
		\bibitem{BertiHassMasm}
		 Berti, M.;  Hassainia, Z.; Masmoudi, N.:
		{\em Time quasi-periodic vortex patches of Euler equation in the plane}.
		Invent. Math. 233, 1279--1391, (2023).
		
%		
%		\bibitem{BKMdnls}
%		Berti M.; Kappeler T.; Montalto R.:
%		\emph{Large KAM tori for perturbations of the defocusing NLS equation}.
%		Astérisque 403, (2018).
		
		%		
		%		\bibitem{BKM} Berti M., T. Kappeler, Montalto R., 
		%		\emph{Large KAM tori for quasi-linear perturbations of KdV}, 
		%		Archive for Rational Mechanics,
		%		https://doi.org/10.1007/s00205-020-01596-2, 2021. 
		
		\bibitem{BM} 
		Berti, M.; Montalto, R.: 
		\emph{Quasi-periodic standing wave solutions of gravity-capillary water waves}.
		MEMO,
		Volume 263, 1273, Memoires AMS, ISSN 0065-9266, (2020).
%		
%		\bibitem{BCzD}
%		Bianchini, R.; Coti Zelati, M.; Dolce, M.:
%		\emph{Linear inviscid damping for shear flows near Couette in the 2D stably stratified regime}.
%		arXiv preprint, arXiv:2005.09058, (2020).
%		
%		\bibitem{Bou}
%		Bourgain, F.:
%		\emph{Quasi-periodic solutions of Hamiltonian perturbations of 2D linear Schrodinger equations}.
%		Ann. of Math. 148, 363--439, (1998).
%		
\bibitem{MaProBia}
		 Biasco, L.; Massetti, J.E.; Procesi, M.:
		\emph{Almost periodic invariant tori for the NLS on the circle}.
		 Ann. Inst. H. Poincar\'e Anal. Non Lin\'eaire 38(3), 711--758, (2021).
		 
		 \bibitem{BMP2} Biasco, L.; Massetti, J.E.; Procesi, M.:
		\emph{Small amplitude weak almost periodic solutions for the 1D NLS}.
		 Duke Math. Journal 172(14), pp. 2643–2714 (2023).

		\bibitem{Bou04}
		Bourgain, J.:
		\emph{On invariant tori of full dimension for 1D
			periodic NLS}.
		J. Funct. Anal. 229, 62--94, (2004).



%		\bibitem{BD}
%		Bridges, T.J.; Dias, F.:
%		\emph{Spatially quasi-periodic gravity-capillary waves}.
%		Contemporary Mathematics 200, 31--46, (1996).
%		
%		\bibitem{BR}
%		Bridges, T.J.; Rowlands, G.:
%		\emph{Instability of spatially quasi-periodic states of the Ginzburg-Landau equation}.
%		Proc. R. Soc. Lond. A 444, 347--362, (1994).
		
		%\bibitem{CCL}
		%Castro, A.; C\'ordoba, D.; Lear, D.:
		%	\emph{On the asymptotic stability of stratified solutions for the 2D Boussinesq equations with a velocity damping term}.
		%	Math. Mod. Meth. Appl. Sc. 29(7), 1227--1277, (2019).
		
		\bibitem{CL21}
		Castro, A.; Lear, D.:
		\emph{Traveling waves near Couette flow for the 2D Euler equation}.
		arXiv preprint, arxiv:2111.03529, (2021).
		
		\bibitem{CL23}
		Castro, A.; Lear, D.:
		\emph{Time periodic solutions for 2D Euler near Taylor-Couette flow}.
		arXiv preprint,   arXiv:2311.06885, (2023).
		
		\bibitem{CLV}
		Constantin, P; La, J.; Vicol, V: 
		\emph{Remarks on a paper by Gravilov: Grad-Shafranov equations, steady solutions of the three dimensional incompressible Euler equations with compactly supported velocities, and applications.}
		 Geom. Funct. Anal. 29, 1773--1793, (2019).
		 
		 \bibitem{CGP} Corsi, L.; Gentile, G.; Procesi, M. \emph{Almost-periodic solutions to the NLS equation with smooth convolution potentials.} Preprint arXiv: 2309.1427 (2023). 
		 
		 \bibitem{CoMoPro}
		 Corsi, L.; Montalto, R.; Procesi, M.:
		 \emph{Almost-periodic response solutions for a forced quasi-linear Airy equation}.
		 J. Dyn. Diff. Eq.  33, 1231--1267, (2021).
		
		\bibitem{CzEW}
		Coti Zelati, M.; Elgindi, T.; Widmayer, K.:
		\emph{Stationary structures near the Kolmogorov and Poiseuille flows in the 2d Euler equations}.
		Arch. Rat. Mech. Anal. 247:12, (2023).
%		arXiv preprints, arXiv:2007.11547, (2020).
		
		%        \bibitem{CM1}
		%       Coughlin, K.T.; Marcus, P.S.:
		%      \emph{Modulated waves in Taylor-Couette flows Part 1. Analysis}.
		%     J. Fluid Mech. 234, 1--18, (1992).
		
		%        \bibitem{CM2}
		%       Coughlin, K.T.; Marcus, P.S.:
		%      \emph{Modulated waves in Taylor-Couette flows Part 2. Numerical simulation}.
		%     J. Fluid Mech. 234, 19--46, (1992).
		
%		\bibitem{CW}
%		Craig, W.; Wayne, C.E.:
%		\emph{Newton’s method and periodic solutions of nonlinear wave equations}.
%		Comm. Pure Appl. Math. 46(11), 1409--1498, (1993).
		
		\bibitem{CF}
		Crouseilles, N.; Faou, E.:
		\emph{Quasi-periodic solutions of the 2D Euler equation}.
		Asymptotic Analysis 81(1), 31--34, (2013).
		
%		\bibitem{DeMa}
%		Deng, Y.; Masmoudi, N.:
%		\emph{Long time instability of the Couette flow in low Gevrey spaces}.
%		To appear in Comm. Pure Appl. Math..

		
		\bibitem{EPsTdL}
		Enciso, A.; Peralta-Salas, D.; Torres de Lizaur, F.:
		\emph{Quasi-periodic solutions to the incompressible Euler equations in dimensions two and higher}.
		Journal Diff. Eq. 354, 170--182, (2023).
		
%		\bibitem{fejoz}
%		F\'ejoz, J.:
%		\emph{D\'emonstration du th\'eor\`eme d’Arnold sur la stabilit\`e du syst\'eme plan\`etaire (d'apr\'es Herman)}.
%		Ergod. Th.  Dynam. Sys.  24, 1521--1582, (2004).
		
		\bibitem{FeoGiu}
		Feola, R.; Giuliani, F.: 
		\emph{Quasi-periodic traveling waves on an infinitely deep fluid under gravity}. 
		To appear on Memoirs AMS (2022).
		
		\bibitem{FrMo}
		Franzoi, L.; Montalto, R.:
		\emph{A KAM approach to the inviscid limit for the 2D Navier-Stokes equations}. To appear on Annales Herni Poincar\'e.
%		arXiv preprint, 
		arXiv:2207.11008, (2022).
		
		\bibitem{FrMaMo}
		Franzoi, L.; Masmoudi, N.; Montalto, R.:
		\emph{Space quasi-periodic steady Euler flows close to the inviscid Couette flow}.
		arXiv preprint, arXiv:2303.03302, (2023).
		
		\bibitem{gravilov}
		Gravilov, A.V.: 
		\emph{A steady Euler flow with compact support.}
		Geom. Funct. Anal. 29, 190--197, (2019).
		
		
		\bibitem{GomIonePa}
			G\'omes-Serrano, J.; Ionescu, A.D.; Park, J.:
		\emph{Quasiperiodic solutions of the generalized SQG equation}.
		arXiv preprint, arXiv:2303.03992.
%		\bibitem{GroWah}
%		Groves, M.D.; Wahl\'el, E.:
%		\emph{Spatial Dynamics Methods for Solitary Gravity-Capillary Water Waves with an Arbitrary Distribution of Vorticity}.
%		SIAM J. Math. Anal. 39(3), 932--964, (2007).
%		
		\bibitem{HaHmMa}
		Hassainia, Z.,; Hmidi, T.; Masmoudi, N.: 
		\emph{KAM theory for active scalar equations}.
		arXiv preprint, arXiv:2110.08615, (2021).
%		
		\bibitem{HaHmRou}
		Hassainia, Z.,; Hmidi, T.; Roulley, E.: 
		\emph{Invariant KAM tori around annular vortex patches for 2D Euler equations}.
		arXiv preprint, arXiv:2302.01311, (2023).
%		
		\bibitem{HaRo}
		Hassainia, Z.; Roulley, E.:
		\emph{Boundary effects on the emergence of quasi-periodic solutions for Euler equations}.
		arXiv preprint, arXiv:2202.10053, (2022).
		
		\bibitem{HmRo}
		Hmidi, T.; Roulley, E.: 
		\emph{Time quasi-periodic vortex patches for quasi-geostrophic shallow water equations}.
		arXiv preprint, arXiv:2110.13751, (2021).
%		
%		
%		%		\bibitem{geng1}
%		%		Geng J., You J.,
%		%		\newblock{\em  A KAM theorem for one dimensional Schr\"odinger equation with
%			%			periodic boundary conditions,}
%		%		\newblock{ J. Diff.  Eq.  209(1)},  1--56, 2005.
%		
%		
%		%		\bibitem{geng2}
%		%		Geng J., You J., 
%		%		\newblock{\em A KAM theorem for Hamiltonian partial differential equations in
%			%			higher dimensional spaces,}
%		%		\newblock{ Comm. Math. Phys.},  262, 343--372, 2006.
%		
%		
%		
%		
%		%		\bibitem{IP-SW1}
%		%		Iooss G., Plotnikov P.,
%		%		\newblock {\it Multimodal standing gravity waves: a completely resonant system}, 
%		%		\newblock J. Math. Fluid Mech., 7(suppl. 1) : S110-S126, 2005.
%		%		
%		%		%\bibitem{IP-SW2}
%		%		%Iooss G., Plotnikov P.,
%		%		%\newblock {\it Existence of multimodal standing gravity waves}, 
%		%		%\newblock { J. Math. Fluid Mech.}, 7(suppl. 3):S349-S364, 2005.
%		%		
%		%		
%		%		\bibitem{IP-Mem-2009}
%		%		Iooss G., Plotnikov P.,
%		%		\newblock {\it Small divisor problem in the theory of three-dimensional water
%			%			gravity waves}, 
%		%		\newblock { Mem. Amer. Math. Soc.}, 200(940):viii+128, 2009.
%		%		
%		%		\bibitem{IP2}
%		%		Iooss G., Plotnikov P.,
%		%		\newblock {\it Asymmetrical tridimensional traveling gravity waves}, 
%		%		\newblock { Arch. Rat. Mech. Anal.}, 200(3):789--880, 2011.
%		
%		\bibitem{IoJi}
%		Ionescu, A.D.; Jia, H.:
%		\emph{Inviscid damping near the Couette flow in a channel}.
%		Comm. Math. Phys. 374, 2015--2096, (2020).
%		
%		\bibitem{IoLos}
%		Iooss, G.; Los, J.:
%		\emph{Bifurcation of spatially quasi-periodic solutions in hydrodynamic stability problems}.
%		Nonlinearity 3, 851--871, (1990).
%		
%		\bibitem{IoMie}
%		Iooss, G.; Mielke, A.:
%		\emph{Bifurcating time-periodic solutions of Navier-Stokes equations in infinite cylinders}.
%		J. Nonlin. Sci. 1, 107--146, (1991).
%		
%		\bibitem{IPT}
%		Iooss, G.; Plotnikov, P.; Toland, J.:
%		\newblock {\it Standing waves on an infinitely deep perfect fluid under gravity}. 
%		\newblock  Arch. Rat. Mech. Anal. 177(3), 367--478, (2005).
%		
%		\bibitem{Kelvin}
%		Lord Kelvin: 
%		\emph{Stability of fluid motion: rectilinear motion of viscous fluid between two parallel plates}. 
%		Phil. Mag. 24(5), 188--196, (1887).
%		
%		\bibitem{Kuksin}
%		Kuksin, S.: 
%		\emph{Hamiltonian perturbations of infinite-dimensional linear systems with an imaginary spectrum}. 
%		Funct. Anal. Appl. 21, 192--205, (1987).
%		
%		\bibitem{LiLin}
%		Li, C.; Lin, Z.:
%		\emph{A resolution of the Sommerfeld paradox}.
%		SIAM J. Math. Anal. 43(4), 1923--1954, (2011).
%		
%		\bibitem{LMZ}
%		Li, H.; Masmoudi, N.; Zhao, W.:
%		\emph{New energy method in the study of the instability near Couette flow}.
%		arXiv preprint, arXiv:2203.10894, (2022).
%		
		\bibitem{LZ}
		Lin, Z.; Zeng, C.: 
		\newblock {\it Inviscid dynamical structures near Couette flow}.
		\newblock Arch. Rat. Mech. Anal. 200, 1075-1097, (2011).
%		
%		\bibitem{LZ13}
%		Lin, Z.; Zeng, C.: 
%		\newblock {\it Unstable manifolds of Euler equations},
%		\newblock Comm. Pure Appl. Math. 46, 1803--1836, (2013).


		
%		
%		%	\bibitem{Mont}
%		%	Montalto, R.:
%		%	\emph{The Navier–Stokes Equation with Time Quasi-Periodic External Force: Existence and Stability of Quasi-Periodic Solutions}.
%		%    J. Dyn. Diff. Eq., 33(3), 1341--1362, (2021).




\bibitem{MoPr021}
Montalto, R.; Procesi, M.:
\emph{Linear Schr\"odinger equation with an almost periodic potential}.
SIAM J. Math. Anal. 53(1), 386--434, (2021).


%		
%		\bibitem{MouVil}
%		Mouhot, C.; Villani, C.:
%		\emph{On Landau damping}.
%		Acta Math. 207(1), 29--201, (2011).
%		
%		\bibitem{Orr}
%		Orr, W.: 
%		\emph{The stability or instability of the steady motions of a perfect liquid and of a viscous liquid}.
%		Proc. Royal Irish Acad. Sec. A: Math. and
%		Phys. Sci. 27, 9--138, (1907-1909).
%		
%		\bibitem{PlTo}
%		Plotnikov, P.;  Toland, J.:
%		\newblock {\it Nash-{M}oser theory for standing water waves}.
%		\newblock { Arch. Rat. Mech. Anal.} 159(1), 1--83, (2001).
%		
%		\bibitem{PV}
%		Pol\'a\v{c}ik, P.; Valdebenito, D.A.:
%		\emph{Existence of quasiperiodic solutions of elliptic equations on $\R^{N+1}$ via center manifold and KAM theorems}.
%		J. Diff. Eq. 262, 6109--6164, (2017).
%		

		\bibitem{PoeschelAP}
		P\"oschel, J.:
		\emph{On the construction of almost periodic solutions
			for a nonlinear Schrodinger equation}.
			Ergod. Th. Dynam. Sys. 22, 1537--1549, (2002).


		\bibitem{Roulley}
		Roulley, E.:
		\emph{Periodic and quasi-periodic Euler-$\alpha$ flows close to Rankine vortices}.
		arXiv preprint, arXiv:2208:13109, (2022).
		
%		\bibitem{Russ}
%		R\"ussmann H., {\it Invariant tori in non-degenerate nearly integrable Hamiltonian systems}. {Regul. Chaotic Dyn.} 6(2), 119-204, (2001).
%		
%		\bibitem{Scheurle}
%		Scheurle, J.:
%		\emph{Quasiperiodic solutions of a semilinear equation in a two-dimensional strip}.
%		Dynamical Problems in Mathematical Physics, Band 26, B. Brosowski and E. Martensen (eds.), 201--223, Frankfurt Am Maim: Lang, 1983.
%		
%		\bibitem{teschl}
%		Teschl G., 
%		{\it Mathematical methods in quantum mechanics. With applications to Schrödinger operators.} 
%		Graduate Studies in Mathematics, 99, American Mathematical Society, Providence, RI, (2009).
%		
%		\bibitem{Valls}
%		Valls, C.:
%		\emph{Existence of quasi-periodic solutions for elliptic equations on a cylindrical domain}.
%		Comment. Math. Helv. 81, 783--800, (2006).
%		
%		\bibitem{Wayne}
%		Wayne, C.E.: 
%		\emph{Periodic and quasi-periodic solutions of nonlinear wave equations via KAM theory}. 
%		Comm. Math. Phys. 127(3):479--528, (1990).
%		
%		\bibitem{YangLin}
%		Yang, J.; Lin, Z.:
%		\emph{Linear inviscid damping for Couette flow in stratified fluid}.
%		J. Math. Fluid Mech. 20, 445--472, (2018).
%		
%		\bibitem{Zill}
%		Zillinger, C.:
%		\emph{Linear inviscid damping for monotone shear flows}.
%		Trans. Amer. Math. Soc. 369(12), 8799--8855, (2017).
		
	\end{thebibliography}

\documentclass[12pt,a4paper,reqno]{article}

\usepackage[latin,english]{babel}
\usepackage[utf8]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{eufrak, color, mathrsfs,dsfont}
\usepackage{times}
\usepackage{geometry}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{tikz}
\usepackage{mathtools}
\usepackage{mathabx}
\usepackage{cases}
\usepackage{braket}
\usepackage[toc,page]{appendix}
\usepackage{pdfsync}
\usepackage{hyperref}
\usepackage{psfrag}
\mathtoolsset{showonlyrefs}
\usepackage[mathscr]{euscript}
\usepackage{math}
\numberwithin{equation}{section}

\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}
\newtheorem{defn}[thm]{Definition}

\theoremstyle{remark}
\newtheorem{rem}[thm]{Remark}

\DeclareMathOperator{\spt}{spt}

\newcommand{\pa}{\partial}

\title{Time almost-periodic solutions of the incompressible Euler equations}

\author{Luca Franzoi, Riccardo Montalto}

\date{}

\begin{document}

\maketitle

%\date{\today}

\begin{abstract}
We construct time almost-periodic solutions (global in time) with finite regularity to the incompressible Euler equations on the torus $\T^d$, with $d=3$ and $d\in\N$ even.
\end{abstract}

\noindent
{\em Keywords:} Fluid dynamics, Euler equations, almost-periodic solutions.
% Nash-Moser theory, 
  %  invariant tori.

\noindent
{\em MSC 2020:}  35Q31, 37N10, 76M45.

%\tableofcontents

\section{Introduction}

The goal of this paper is to construct \emph{time almost-periodic} solutions (infinite dimensional invariant tori) of the Euler equations
\begin{equation}\label{euler.eq}
\begin{aligned}
& 	\pa_{t} u + u\cdot \nabla u + \nabla p = 0 \,, \quad {\rm div} \,u = 0\,, \\
& u : \R \times \T^d \to \R^d\,, \quad p : \R \times \T^d \to \R\,,
	\end{aligned}
\end{equation}
on the $d$-dimensional torus $\T^d$, $\T:= \R / 2\pi\Z$, where either $d=3$ or $d\geq 2$ is any even positive integer. 
These solutions extend the works ot Crouseille \& Faou \cite{CF} (in dimension 2) and Enciso, Peralta Salas \& Torres de Lizaur \cite{EPsTdL} (in dimensions 3 or even) from time quasi-periodic to time almost-periodic solutions.
In fact, the construction here follows closely the one in \cite{EPsTdL}.

We need to specify how a smooth solution of the Euler equations \eqref{euler.eq} is called \emph{almost-periodic} in this paper. We need some preliminaries.

Let $\cC_{\rm div}^{s}(\T^d,\R^d)$, with $s\in\N \cup\{+\infty\}$, be the space of $\cC^s$-smooth, divergence free $d$-dimensional vector fields on $\T^d$. This space is a Banach space if $s< \infty$ and a Fr\'echet space when $s=\infty$. We endow it with the system of seminorms $(\|\,\cdot\,\|_{n,\infty})_{n\in\{0,1,...,s\}}$ defined by 
\begin{equation}
	 \|f\|_{n,\infty} := \sup_{x\in \T^d} \max_{\begin{subarray}{c}
	 \alpha\in \N_0^d \\
	  |\alpha|=n
	  \end{subarray}} |\pa_{x}^\alpha f(x) | \,, \quad n=0,1,...,s\,;
\end{equation}
throughout the paper, for sake of simplicity in the notation, $|\,\cdot \,|$ denotes the standard Euclidean norm, without specifying the dimension of the evaluated object, which will be clear from the context each time. We denote by $\ell^\infty(\N,\N)$ the set of sequences in $\N$ that are bounded.
 Let $(J_k)_{k\in\N}\in \ell^\infty(\N,\N)\setminus\{0\}$ be given and, for a fixed $m\in\{1,...,d-1\}$, we define the sequence
\begin{equation}\label{sequence.Nk}
	(N_k)_{k\in\N} \in \ell^\infty(\N,\N)\,, \quad N_k:=(d-m) J_{k} \in \N \,, \quad k\in\N\,.
\end{equation}
We define the infinite dimensional torus $(\T^{N_k})_{k\in\N}$ and its "tangent space" $(\R^{N_k})_{k\in\N}$ as 
\begin{equation}
\begin{aligned}
		& (\T^{N_k})_{k\in\N} := \big\{ \theta = (\theta_{k})_{k\in\N} \, :\, \theta_{k} \in \T^{N_k} , \ | \theta|_{\infty} < \infty  \big\}\,, \\
		& 	(\R^{N_k})_{k\in\N} := \big\{ \nu = (\nu_{k})_{k\in\N} \, :\, \nu_{k} \in \R^{N_k}  , \ | \nu |_{\infty} < \infty   \big\}\,,
\end{aligned}
\end{equation}
where we defined $| \nu |_{\infty}:= \sup_{k\in\N} |\nu_{k}|$. Note that, since $(N_{k})_{k\in\N}$ is bounded, then $| \theta |_{\infty}\leq (2\pi)^{\| (N_k)\|_{\ell^\infty}} < \infty$ for any sequence $\theta = (\theta_{k})_{k\in\N}$.

\begin{defn}
	Let $s\in\N \cup \{+\infty\}$. We say that $u(t,x)$ is \emph{time almost-periodic}
	if there exists a sequence  of vectors $\nu \in (\R^{N_k})_{k\in\N} $ and a $\cC^1$-smooth embedding $U:(\T^{N_k})_{k\in\N}\to \cC_{\rm div}^{s} (\T^d,\R^d)$ such that the velocity field $u(t,x)$ can be written as
	\begin{equation}\label{u(t).as.embedding}
		u(t,\,\cdot\,) = U(\vartheta)|_{\vartheta=\theta+\nu t} \,, \quad \text{ for some} \quad \theta\in (\T^{N_k})_{k\in\N} \,,
	\end{equation}
	and the  sequence of frequency vectors $\nu=(\nu_{k})_{k\in\N}$ is \emph{non-resonant}, meaning that
%	% for any $N\in\N$ and any $K\subset \N$, $|K|=N$, we have that $\sum_{k\in K} \nu_{k} \ell_{k}\neq 0 $ for any $(\ell_{k})_{k\in K} \in \Z^N$.
%	\begin{equation}\label{nonres}
%		\sum_{k\in \N} \nu_{k} \cdot \ell_{k}\neq 0 \quad \ \forall\, \ell_{k} \in \Z^{N_k}, \ k \in \N, \ \text{ with } \ \ (\ell_{k})_{k\in \N} \neq (0)_{k\in\N}  \,.
%	\end{equation}
\begin{equation}\label{nonres}
	\sum_{k\in \N} \nu_{k} \cdot \ell_{k} \neq 0\,, \quad \forall\,\ell=(\ell_{k})_{k\in\N} \in (\Z^{N_k})_{k\in\N} \ \ \text{ with } \ \ 0< |\ell|_{\eta} <\infty  \,,
\end{equation}
where, for a fixed $\eta>0$, we define $|\ell|_{\eta} := \sum_{k\in \N} k^\eta |\ell_{k}| $.
%\begin{equation}
%	|\ell|_{\eta} := \sum_{k\in \N} k^\eta |\ell_{k}| \,.
%\end{equation}
Note that $|\ell|_{\eta}< \infty$ implies that $\ell_{k}\neq 0 \in \Z^{N_k}$ only for finitely many $k\in\N$.
\end{defn}

\begin{defn}\label{C1-smooth.def}
	By saying that the map  $U:(\T^{N_k})_{k\in\N}\to \cC_{\rm div}^{s} (\T^d,\R^d)$ is $\cC^1_b$ (${\mathcal C}^1$ and bounded), we mean that, $U$ is a Frechet-differentiable map with continuous Frechet derivative and for any $n\in \{ 0, 1, \ldots, s\}$, there exists a constant $C_n>0$ such that
%	\begin{equation}
%	\begin{aligned}
%		& \sup_{\vartheta \in (\T^{N_k})_{k\in\N}} \|  U(\vartheta) \|_{n,\infty} \leq C_n \,, \\
%		&	\sup_{\vartheta \in (\T^{N_k})_{k\in\N}} \| \di_{\vartheta} U(\vartheta) [\nu] \|_{n,\infty} \leq C_n | \nu |_{\infty} \quad \forall\,\nu \in (\R^{N_k})_{k\in\N}\,, 
%	\end{aligned}
%	\end{equation}
	\begin{equation}
	\begin{aligned}
		& \sup_{\vartheta \in (\T^{N_k})_{k\in\N}} \|  U(\vartheta) \|_{n,\infty} \leq C_n   \,,  \\
		&	\sup_{\vartheta \in (\T^{N_k})_{k\in\N}} \| \di_{\vartheta} U(\vartheta) [\wh\vartheta] \|_{n,\infty} \leq C_n | \wh\vartheta |_{\infty} \quad \forall\, \wh\vartheta \in (\R^{N_k})_{k\in\N}\,, 
	\end{aligned}
\end{equation}
	where the linear operator $\di_{\vartheta} U(\vartheta) : ( \R^{N_k})_{k\in\N} \to \cC_{\rm div}^{s} (\T^d,\R^d)$ is the Fr\'echet differential of the embedding $U(\vartheta)$ at the point $\vartheta\in (\T^{N_k})_{k\in\N}$. 
\end{defn}
With this definition of a $\cC^1_{b}$ embedding, we have that the function $u(t,\,\cdot\,)$ is $\cC^1$ with respect to $t\in\R$, by \eqref{u(t).as.embedding} and $\pa_{t} u(t,\,\cdot\,) = \di_{\vartheta} U(\theta + \nu t) [\nu] \in \cC_{\rm div}^{s}(\T^d,\R^d)$.

We will look for solutions where the embedding $U$ is \emph{non-symmetric}, or \emph{non-traveling}, in the sense that, for any $\vartheta\in(\T^{N_k})_{k\in\N}$, the divergence-free vector field $U(\vartheta)$ is not invariant under any 1-parameter group of translations on $\T^d$. In this way, we ensure that the solution $u(t,x)$ depends effectively on all $d$ coordinates and we do not have any reduction to solutions of lower dimensions by traveling directions.

The statement of the main result is the following.

\begin{thm}\label{teorema}
	{\bf (Time almost-periodic solutions of the Euler equations).}
	Assume that the dimension $d$ is either 3 or even. Let $S\in \N$ be fixed. There exists $\varepsilon_{0}\in (0,1)$ small enough such that, for any $\varepsilon\in (0,\varepsilon_{0})$ and for any sequence of frequencies $\nu \in (\R^{N_k})_{k\in\N}\setminus\{0\}$ satisfying
	\begin{equation}\label{small.freq}
		\sup_{k\in \N} \varepsilon^{-(S + 1)(k-1)} |\nu_{k}| <\infty \,,
	\end{equation}
	there exists a non-symmetric $\cC^{1}_{b}$ embedding $U:(\T^{N_k})_{k\in\N}\to \cC_{\rm div}^S(\T^d,\R^d)$ and a family of initial data $u_{\theta}\in \cC_{\rm div}^S(\T^d,\R^d)$, $\theta\in (\T^{N_k})_{k\in\N}$, such that $u(t,\,\cdot\,)= U(\theta+\nu t)$, with $u(0,\,\cdot\,)=u_{\theta}$, is a solution of \eqref{euler.eq} with pressure $p(t,\,\cdot\,)= P(\theta+\nu t)$, where
	\begin{equation}
		P(\vartheta) :=(-\Delta)^{-1}\big[ {\rm div} (U(\vartheta)\cdot \nabla U(\vartheta)  )\big] : (\T^{N_k})_{k\in\N} \to \cC^S(\T^d)\,.
 	\end{equation}
	As a consequence, if the sequence $\nu=(\nu_{k})_{k\in\N}$ is non-resonant, namely it satisfies \eqref{nonres}, then the solution $u(t,x)$ is time almost-periodic.  
\end{thm}

\begin{rem}
	As it will be clear from the construction in the following section, the embedding $U(\vartheta)$ is determined as a combination of infinitely many embedding $U_k(\vartheta_{k})$, with $\vartheta_{k}\in \T^{N_k}$, which coincides with the embedding constructed in \cite{EPsTdL}, with the size of the embedding $U_{k}$ becoming smaller and smaller  as $k\to \infty$. The major difference in the analysis with respect to \cite{EPsTdL} is that we have to effectively prove the smoothness of the embedding and the regularity of the vector field. This is not trivial. 
\end{rem}

\begin{rem}
	The condition \eqref{nonres} of irrationality for the sequence of frequencies $\nu\in(\R^{N_k})_{k\in\N}\setminus \{0\}$ is not necessary in the construction of the embedding $U$. Depending on relations between all the frequencies, we may obtain embedding for lower dimensional tori, either finite dimensional (quasi-periodic or periodic) or still infinite dimensional (that is, almost-periodic). On the other hand, the control on the frequency vectors in \eqref{small.freq} is required to ensure that the solution $u(t,x)$ is indeed a finitely smooth vector field and a simpler control on the norm $|\nu|_\infty$ is not enough. At the physical level, is also implies that we obtain solutions whose leading order frequencies of oscillations are only finitely many and the almost-periodicity in time is due to the presence of infinitely oscillations with smaller and smaller frequencies.
%	The condition can be even restricted to obtain an infinitely smooth solution.
\end{rem}

\paragraph{Related results.} 
In the last years, there has been a discrete surge of works proving the existence of time quasi-periodic waves for PDEs arising in fluid dynamics.
With the exception of the aforementioned works \cite{CF} and \cite{EPsTdL}, there type of results in literature are proved by means of KAM for PDEs techniques, to deal with the presence of small divisors issues and consequent losses of regularity.
For the two dimensional water waves equations, we mention
%Iooss, Plotnikov & Toland [27] for periodic standing waves, 
Berti \& Montalto \cite{BM}, Baldi, Berti, Haus \& Montalto \cite{BBHM} for time quasi-periodic standing waves and Berti, Franzoi \& Maspero \cite{BFM},
\cite{BFM21}, Feola \& Giuliani \cite{FeoGiu} for time quasi-periodic traveling wave solutions. 
Recently, the existence of time quasi-periodic solutions was proved for the contour dynamics of vortex patches in active scalar equations. 
We mention Berti, Hassainia \& Masmoudi \cite{BertiHassMasm} for vortex patches of
the Euler equations close to Kirchhoff ellipses, Hmidi \& Roulley \cite{HmRo} for the 
quasi-geostrophic shallow water equations,  Hassainia, Hmidi \& Masmoudi \cite{HaHmMa} for generalized surface quasi-geostrophic equations, Roulley \cite{Roulley} for Euler-$\alpha$ flows, Hassainia \& Roulley \cite{HaRo} for Euler equations in the unit disk close to Rankine vortices and Hassainia, Hmidi \& Roulley \cite{HaHmRou} for 2D Euler annular vortex patches.
%We mention Berti, Hassainia \& Masmoudi \cite{BertiHassMasm} for vortex patches of
%the Euler equations close to Kirchhoff ellipses, Hmidi \& Roulley \cite{HmRo} for the 
%quasi-geostrophic shallow water equations,  Hassainia, Hmidi \& Masmoudi \cite{HaHmMa} for generalized surface quasi-geostrophic equations, Roulley \cite{Roulley} for Euler-$\alpha$ flows, Hassainia \& Roulley \cite{HaRo} for Euler equations in the unit disk close to Rankine vortices and Hassainia, Hmidi \& Roulley \cite{HaHmRou} for 2D Euler annular vortex patches.
%All the aforementioned results concern 2D Euler equations.
Time quasi-periodic solutions were also constructed for the 3D Euler equations with time quasi-periodic external force \cite{BM20} and for the forced 2D Navier-Stokes
equations \cite{FrMo} approaching in the zero viscosity limit time quasi-periodic solutions of the 2D Euler equations for all times. 

The existence of other non-trivial invariant structures is also a topic of interest in  fluid dynamics. In particular, for the Euler equations in two dimension close to shear flows, we mention the works by Lin \& Zeng \cite{LZ} and Castro \& Lear \cite{CL21} for periodic traveling waves close the Couette flow,  by Coti Zelati, Elgindi \& Widmayer \cite{CzEW} for stationary waves around non-monotone shears, by  Franzoi, Masmoudi \& Montalto \cite{FrMaMo} for quasi-periodic traveling waves close to the Couette flow, and the recent work by Castro \& Lear \cite{CL23} for time periodic rotating solutions close to the Taylor-Couette flow.

Concerning the existence of almost periodic solutions by means of KAM methods, we mention P\"oschel \cite{PoeschelAP}, Bourgain \cite{Bou04}, Biasco-Massetti-Procesi \cite{MaProBia}, \cite{BMP2} and Corsi-Gentile-Procesi \cite{CGP}. In all these results the authors consider semilinear NLS type equations with external parameters. For PDEs with unbounded perturbations (with external parameters as well) we mention Montalto-Procesi \cite{MoPr021} and Corsi-Montalto-Procesi \cite{CoMoPro}.

\noindent
We remark that our result is the first one concerning existence of almost-periodic solutions for an autonomous quasi-linear PDEs in higher space dimension and it is obtained with non-KAM techniques. 
%We finally mention that time quasi-periodic solutions for the Euler equations were constructed also by Crouseilles \& Faou \cite{CF} in 2D, with a very recent extension by Enciso, Peralta-Salas \& de Lizaur \cite{EPsTdL} in 3D and even dimensions: we remark that these latter solutions are engineered so that there are no small divisors issues to deal with, with consequently much easier proofs and a drawback of not having information on the eventual stability of the solutions. 

\bigskip

\medskip

%	\paragraph{Acknowledgments.} 
\noindent
{\bf Acknowledgments. }
	The work of the authors Luca Franzoi and Riccardo Montalto is funded by the European Union, ERC STARTING GRANT 2021, "Hamiltonian Dynamics, Normal Forms and Water Waves" (HamDyWWa), Project Number: 101039762. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. 

\noindent
The work of the author Riccardo Montalto is also supported by PRIN 2022 "Turbulent effects vs Stability in Equations from Oceanography", project number: 2022HSSYPN. 

\noindent
Riccardo Montalto is also supported by INDAM-GNFM. 

\bigskip
\medskip

%\paragraph{Notations.} 
\noindent
{\bf Notations.}
In this paper, we use the following notations:
%\\[1mm]
%\noindent $\bullet$ $\N := \{1,2,...,n,...\}$ and $\N_0:= \{0\}\cup \N$;
%\\[1mm]
%\noindent $\bullet$ $|\,\cdot\,|$ denotes the Euclidean norm in the dimension of the evaluated object, without any further specification;
\\[1mm]
\noindent $\bullet$ $B_{d,\rho}(\tp) := \{ x\in\R^d \,:\, |x-\tp| < \rho \}$, with $\tp \in \R^d$ and $\rho >0$;
%denoted the open ball in $\R^d$ (or eventually $\T^d$) centered at $\tp \in \R^d$ with radius $\rho>0$.
%\\[1mm]
%\noindent $\bullet$ For any $n\in\N_0$, we define the seminorm $ |f|_{n,\infty} := \sup_{x\in\spt(f)} \max_{\alpha\in \N_0^d, \, |\alpha|=n} |\pa_{x}^\alpha f(x) |$
%\begin{equation}
%		|f|_{n,\infty} := \sup_{|x|\leq \varepsilon^{k}} \max_{\alpha\in \N_0^d, \, |\alpha|=n} |\pa_{x}^\alpha f(x) |
%\end{equation}
%where $f:\R^{m_1}\to \R^{m_2}$;
\\[1mm]
\noindent $\bullet$ $\cC^{s}(\T^{m_1},\R^{m_2}):= \big\{ f:\R^{m_1}\to \R^{m_2} \,:\, \|f \|_{n,\infty} <\infty \ \ \forall\, n\in \N \cup \{ 0 \}, \, n\leq s  \big\}$, $s \in \N \cup \{\infty\}$;
%\\[1mm]
%\noindent $\bullet$ $\cC^{\infty}(\T^{m_1},\R^{m_2}):= \big\{ f:\R^{m_1}\to \R^{m_2} \,:\, |f|_{n,\infty} <\infty \ \ \forall\, n\in\N_0 \big\}$;
%\\[1mm]
%\noindent $\bullet$ $\cC_{\rm div}^{S}(\T^{d},\R^{d}):= \big\{ f\in \cC^{S}(\T^{d},\R^{d}) \,:\, {\rm div} \, f = 0 \big\} $, $S\in \N \cup \{\infty\}$;
\\[1mm]
\noindent $\bullet$ $\cC^{\infty}(X,\R) := \cC^{\infty}(X)$, with $X=\T^d, \R^d$;
\\[1mm] 
\noindent $\bullet$ $a \lesssim b$ stands for $a \leq C b$, for some constant $C>0$;
\\[1mm] 
\noindent $\bullet$ $a \lesssim_{n} b$ stands for $a \leq C_{n} b$, for some constant $C_n>0$ depending on $n$.

%\section{The construction of the solutions}
\section{Proof of Theorem \ref{teorema}}

The scheme follows essentially the one proposed in \cite{EPsTdL}, with the required adaptations. The key starting point is the existence of smooth, compactly supported stationary solutions of the Euler equations. In $d=3$, this is celebrated result by Gravilov \cite{gravilov} (see also \cite{CLV}), whereas in even dimension it has been proved in \cite{EPsTdL}. We recall the statement of the result of the latter.

\begin{prop}\label{gravi.sol}
	{\bf (Smooth stationary Euler flows with compact support - Proposition 2, \cite{EPsTdL}).}
	If $d=3$ or $d\in\N$ is even, there exists a smooth, compactly supported solution $v(x)\in \cC_{\rm div}^\infty(\R^d,\R^d) $, with pressure $p_{v}(x)\in \cC^\infty(\R^d)$, of
	\begin{equation}\label{euler.stat}
		v\cdot \nabla v + \nabla p_{v} = 0 \,, \quad {\rm div} \,v = 0\,, \quad x \in \R^d\,.
	\end{equation}
\end{prop}
 Without any loss of generality, we assume that $\spt(v), \spt(p_v) \subseteq B_{d,1}(0) \subset \R^d$. Then, given $S\in\N$  and for any $k\in\N$, we define the rescaled functions, for any $\varepsilon\in (0,1)$ small enough,
 \begin{equation}
 	v_k(x) := \varepsilon^{(S + 1)(k-1)} v(\varepsilon^{-k}x) \,,\quad p_{v_{k}}(x) := \varepsilon^{2(S + 1)(k-1)} p_{v}(\varepsilon^{-k}x) \,.
 \end{equation}
 A straightforward computation shows that $v_{k}(x)\in \cC_{\rm div}^\infty(\R^d,\R^d) $ is also a solution of \eqref{euler.stat} with pressure $p_{v_{k}}(x)\in \cC^\infty(\R^d)$ with compact support
\begin{equation}
	\spt (v_{k}), \spt(p_{v_k}) \subseteq B_{d,\varepsilon^{k}}(0) \subset \R^d\,,
\end{equation}
and we have the control on the seminorms, for any $n\in\N_0$,
%\begin{equation}
%	\sup_{|x|\leq \varepsilon^{k}} \max_{\alpha\in \N_0^d, \, |\alpha|=n} |\pa_{x}^\alpha v_{k}(x) | \leq C_n \varepsilon^{k(S-n)-S} \,, \quad \sup_{|x|\leq \varepsilon^{k}} \max_{\alpha\in \N_0^d, \, |\alpha|=n} |\pa_{x}^\alpha p_{v_{k}}(x) | \leq C_n \varepsilon^{k(2S-n)-2S} 
%\end{equation}
\begin{equation}\label{rescaled.esti}
	\|v_{k}\|_{n,\infty}  \leq C_n \varepsilon^{k(S + 1-n)-S - 1} \,, \quad \|p_{v_{k}}\|_{n,\infty} \leq C_n \varepsilon^{k(2S + 2-n)-2S - 2} \,,
\end{equation}
for some constant $C_n>0$ independent of $\varepsilon \in (0,1)$ and $k\in\N$.
We remark that, as soon as $n> S + 1$, the seminorms $\|v_{k}\|_{n,\infty}$ start to diverge with respect to $k\to\infty$ as $\varepsilon^{-(n-S - 1)k}$ for $\varepsilon\in (0,1)$, whereas the seminorms $\|p_{v_{k}}\|_{n,\infty}$ start to diverge when $n>2S + 2$. 

Moreover, for $\varepsilon\in (0,1)$ small enough and independent of $k\in\N$, we define the periodicized versions
\begin{equation}
	\overline{v_{k}} (x) := \sum_{q\in \Z^d} v_{k}(x+ 2\pi q) \,, \quad 	\overline{p_{v_{k}}} (x) := \sum_{q\in \Z^d} p_{v_{k}}(x+ 2\pi q)\,, \quad x \in \T^d\,.
\end{equation}

%Let now $m\in\{1,...,d-1\}$ and fix a sequence $(J_{k})_{k\in\N} \in \ell^\infty(\N,\N)$ such that $N_k \leq (d-m)J_k$ for any $k\in\N$.
%Let now $m\in\{1,...,d-1\}$ be fixed as in \eqref{sequence.Nk}.
We recall the sequence $(N_k)_{k\in\N}\in\ell^\infty(\N,\N)$ in \eqref{sequence.Nk} is determined by a fixed $m\in \{1,...,d-1\}$ and a fixed sequence $(J_k)_{k\in\N}\in \ell^\infty(\N,\N)$.
The choice of $m \in \{1,...,d-1\}$ induces the splitting $\T^d = \T^m \times \T^{d-m}$ and we write
%\begin{equation}
%	\T^d \ni x = (x',x'')  \in\T^m \times \T^{d-m}\,.
%\end{equation}
\begin{equation}
	\T^d \ni x = (x',x'')  \in\T^m \times \T^{d-m}\,,\quad  \nabla=(\nabla',\nabla'') := (\nabla_{x'},\nabla_{x''}) \,.
\end{equation}
We select a sequence of points $(\ty_{k,j})_{k\in\N, \, j=1,..,J_k}\subset \T^m$ with the properties that:
\\[1mm]
\noindent {\bf (A)} For any $k_1,k_2\in\N$, $j_1=1,...,J_{k_1}$, $j_2=1,...,J_{k_2}$, with $(k_1,j_1)\neq (k_2,j_2)$, we have
\begin{equation}
	\overline{B_{m,2\varepsilon^{k_1}}(\ty_{k_1,j_1})} \cap \overline{B_{m,2\varepsilon^{k_2}}(\ty_{k_2,j_2})} = \emptyset \,;
\end{equation}
%\noindent {\bf (B)} We have $\big|\T^m \setminus \big( \bigcup_{k\in\N} \bigcup_{j=1}^{J_k} \overline{B_{2\varepsilon^{k}}(\ty_{k,j})} \big)\big|\geq \frac12 \tA_{m} >0$, where $\tA_{m} = |\T^m|=(2\pi)^m$.
\noindent {\bf (B)} We have $\big|\T^m \setminus \big( \bigcup_{k\in\N} \bigcup_{j=1}^{J_k} \overline{B_{m,2\varepsilon^{k}}(\ty_{k,j})} \big)\big|\geq \frac23 |\T^m|>0$.
\\[1mm]
\indent The existence of such sequence of points with these desired properties is proved in the following lemma.
%\begin{lem}
%	There exists $\varepsilon_{0}\in (0,1)$ small enough and a choice of distinct points $(\ty_{k,j})_{k\in\N, \, j=1,..,J_k}\subset \T^m$, such that, for $\varepsilon\in (0,\varepsilon_{0})$, the conditions {\bf (A)} and {\bf (B)} are satisfies.
%\end{lem}
\begin{lem}
	There exist $\varepsilon_{0}=\varepsilon_{0}\big(m,\| (J_k) \|_{\ell^\infty}\big)\in (0,1)$ small enough and a choice of infinitely many distinct points $(\ty_{k,j})_{k\in\N, \, j=1,..,J_k}\subset \T^m$, such that the following holds. For $\varepsilon>0$, we define iteratively the sets
	\begin{equation}\label{E.sets}
		E_0:= \emptyset \,, \quad E_k :=  E_{k-1} \cup \bigcup_{j=1}^{J_k}\overline{ B_{m,2\varepsilon^{k}}(\ty_{k,j}) } \,, \quad k \in \N \,.
	\end{equation}
	Then, for any $\varepsilon\in (0,\varepsilon_{0})$ and for any $k\in\N$, we have:
	%	 \begin{equation}
		%	 	E_{k} \subsetneq \T^m \setminus \bigcup_{n=1}^{k} E_{n-1} \,, \quad  	\overline{B_{2\varepsilon^{k}}(\ty_{k_1,j_1})} \cap \overline{B_{2\varepsilon^{k}}(\ty_{k_2,j_2})} = \emptyset \,,\quad  \forall j_1,j_2 =1,...,J_k \,,
		%	 \end{equation}
	%	 and
	%	 \begin{equation}
		%	 	|\T^m \setminus E_{k} | =\Big( 1- \sum_{n=1}^{k} 2^{-n-1}  \Big) |\T^m| \,.
		%	 \end{equation}
	\\[1mm]
	%	 \noindent $(i)$ $ E_{k} \subsetneq \T^m \setminus \bigcup_{n=1}^{k} E_{n-1} $;
	\noindent $(i)$ $E_{k-1} \cap \bigcup_{j=1}^{J_k} \overline{ B_{m,2\varepsilon^{k}}(\ty_{k,j})} =\emptyset$
	\\[1mm]
	\noindent $(ii)$ $ \overline{B_{m,2\varepsilon^{k}}(\ty_{k_1,j_1})} \cap \overline{B_{m,2\varepsilon^{k}}(\ty_{k_2,j_2})} = \emptyset $ for any $j_1,j_2 =1,...,J_k$;
	\\[1mm]
	%	 \noindent $(iii)$ $|\T^m \setminus \bigcup_{n=1}^{k} E_{n} | =\Big( 1- \sum_{n=1}^{k} 2^{-n-1}  \Big) |\T^m| $.
	\noindent $(iii)$ $\T^m \setminus E_k $ is open and $|\T^m \setminus  E_{k} | \geq \Big( 1- \sum_{n=1}^{k} 4^{-n}  \Big) |\T^m| $.
	\\[1mm]
	As a consequence, conditions {\bf (A)} and {\bf (B)} are satisfied.
\end{lem}
%\begin{proof}
%	Pick any choice of distinct points $\ty_{1,1},...,\ty_{1,J_1} \in \T^m$ and we define the set $E_1 := \T^m \setminus \bigcup_{j=1}^{J_1} \overline{B_{2\varepsilon}(\ty_{1,j})}$. It is clear there exists $\varepsilon_{0}= \varepsilon_{0} (m,J_1)\in (0,1)$ small enough such that, for any $\varepsilon \in (0,\varepsilon_{0})$, the closed balls $\big(\overline{B_{2\varepsilon}(\ty_{1,j})}\big)_{j=1,...,J_1}$ are pairwise disjoint  and , 
%	\begin{equation}
	%		|E_1| = (2\pi)^m(1 - C_m J_1 \varepsilon^m ) \geq \tfrac34 (2\pi)^m >0\,,
	%	\end{equation}
%	as soon as $\varepsilon<\varepsilon_{0} \leq  \big( 4C_m J_1 \big)^{-1/m}$.
%	We define inductively the sequence of sets  and $E_{k+1} := E_{k} \setminus\bigcup_{j=1}^{J_{k+1}} \overline{B_{2\varepsilon^{k+1}}(\ty_{k+1,j})}$, for any $k\geq 2$.
%%	\begin{equation}
	%%		E_1 := \T^m \setminus \bigcup_{j=1}^{J_1} \overline{B_{2\varepsilon}(\ty_{1,j})}
	%%	\end{equation}
%\end{proof}
\begin{proof}
	In the following, we use that $|B_{m,2r}(\ty) | = C_{m} r^{m} |\T^m| $, where the explicit constant $C_m := \big( \pi^{m/2} \Gamma(\tfrac{m}{2}+1) \big)^{-1}\in (0,1)$ depends only on the dimension $m\in\N$, where $\Gamma$ is the Euler Gamma function.
	
	We argue by induction. Let $k=1$.
	We pick an arbitrary choice of distinct points $\ty_{1,1},...,\ty_{1,J_1} \in \T^m$ and, for $\varepsilon>0$, we define the set $E_1 := \bigcup_{j=1}^{J_1} \overline{B_{m,2\varepsilon}(\ty_{1,j})}$. By \eqref{E.sets}, item $(i)$ is automatically satisfied for $k=1$. We define $\varepsilon_{1,1}:= \tfrac14 \min \{|\ty_{1,j_1}-\ty_{1,j_2}| \,:\, 1\leq j_1 < j_2\leq J_1  \}$. Then, for any $\varepsilon \in (0,\varepsilon_{1,1})$, the closed balls $\big(\overline{B_{m,2\varepsilon}(\ty_{1,j})}\big)_{j=1,...,J_1}$ are pairwise disjoint, that is, item $(ii)$ is satisfied when $k=1$. Clearly, we also have that $\T^m \setminus E_1$ is open, since $E_1$ is a finite union of closed sets. Moreover, using that $C_m\in(0,1)$, we compute
	\begin{equation}
		|\T^m \setminus E_1| =(1 - C_m J_1 \varepsilon^m ) |\T^m| \geq (1 - J_1 \varepsilon^m ) |\T^m|\geq  \tfrac34 |\T^m| \,,
	\end{equation}
	as soon as $\varepsilon<\varepsilon_{1,2} :=  \big( 4 J_1 \big)^{-1/m}$. Therefore, choosing $\varepsilon_{0} \leq \min\{\varepsilon_{1,1},\varepsilon_{1,2} \}$, we conclude that $(i)$, $(ii)$ and $(iii)$ are satisfied for $k=1$.
	
	We now assume that the claims  $(i)$, $(ii)$ and $(iii)$ are satisfied for some $k\in\N$ and we prove them for $k+1$. We set
	\begin{equation}\label{scelta.eps0}
		\varepsilon_{0} := \min \{ \varepsilon_{1,1} , (4  \| (J_k) \|_{\ell^\infty})^{-1/m} \} \leq \min \{ \varepsilon_{1,1},\varepsilon_{1,2}\} \,.
	\end{equation}
	By \eqref{scelta.eps0}, there exist $J_{k+1}$ distinct points $\ty_{k+1,1},...,\ty_{k+1,J_{k+1}} \in \T^m \setminus E_{k}$,  with  $J_{k+1} \leq \| (J_k) \|_{\ell^\infty}$,  such that, for any $\varepsilon \in (0,\varepsilon_{0})$ we have that the $J_{k+1}$ balls $\overline{B_{m,2\varepsilon^{k+1}}(\ty_{k+1,1})}$,  ..., $\overline{B_{m,2\varepsilon^{k+1}}(\ty_{k+1,J_{k+1}})}$ are contained in $\T^m \setminus E_k$ and they are disjoint, namely they satisfy items $(i)$ and $(ii)$ at the step $k+1$. This follows from the fact that, by the induction assumption on $(iii)$, we have that $\T^m \setminus E_k$ is open with measure $| \T^m \setminus E_k| > \tfrac23 |\T^m|$, whereas the measure of the finite union of closed disjoint balls is estimated, for any $\varepsilon \in (0,\varepsilon_{0})$ with $\varepsilon_{0}$ as in \eqref{scelta.eps0}, by
	\begin{equation}\label{measure.balls}
		\begin{aligned}
			\Big|   \bigcup_{j=1}^{J_{k+1}} \overline{B_{m,2\varepsilon^{k+1}}(\ty_{k+1,j})}\, \Big| & = C_m J_{k+1} \varepsilon^{(k+1)m} |\T^m| \\
			&\leq  \frac{1}{4^{k+1}} \frac{J_{k+1}}{\| (J_k) \|_{\ell^\infty}^{k+1}} |\T^m| \leq  \frac{1}{4^{k+1}} |\T^m| < \frac23 |\T^m| \,,
		\end{aligned}
	\end{equation}
	which implies the existence of the $J_{k+1}$ points $\ty_{k+1,1},...,\ty_{k+1,J_{k+1}}$ in the open and bounded set $\T^m\setminus E_{k}$ with the desired properties. Therefore, let $E_{k+1}$ be defined as in \eqref{E.sets}. Clearly, $E_{k+1}$ is closed, which also implies that $\T^m\setminus E_{k+1}$ is open. By \eqref{measure.balls} and item $(ii)$ at the step $k+1$, we also deduce that,
	\begin{equation}
		|\T^m \setminus E_{k+1} | = |\T^m \setminus E_{k}| - 	\Big|   \bigcup_{j=1}^{J_{k+1}} \overline{B_{m,2\varepsilon^{k+1}}(\ty_{k+1,j})}\, \Big| \geq \Big( 1 - \sum_{n=1}^k 4^{-n} - 4^{-(k+1)} \Big)|\T^m|  \,,
	\end{equation}
	which is indeed the estimate in item $(iii)$ at the step $k+1$. This closes the induction argument and concludes the proof.
	%	\begin{equation}
		%		\Big| (\T^m \setminus E_{k})\setminus \bigcup_{j=1}^{J_{k+1}} \overline{B_{2\varepsilon^{k+1}}(\ty_{k+1,j})}  \Big| = \Big( 1- \sum_{n=1}^{k} 2^{-n-1}   \Big)|\T^m| -  C_m 
		%	\end{equation}
\end{proof}

%As a last preliminary, for any $k\in\N$ we denote $\wt\nu_{k} := \imath_{k}(\nu_{k}) \in\R^{(d-m)J_k}$, where $\imath_{k}:\R^{N_k}\to \R^{(d-m)J_k}$ is the canonical embedding such that, when $N_k=(d-m)J_k$, we have $\imath_{k} = {\rm id}_{\R^{N_k}}$.
As a last preliminary, we take a sequence of frequency vectors $\nu = (\nu_{k})_{k\in\N}\in (\R^{N_k})_{k\in\N}$, where  $\nu_{k} = (\nu_{k,1},...,\nu_{k,J_k}) \in \R^{N_k}$, with $\nu_{k,j}\in\R^{d-m}$, recalling \eqref{sequence.Nk}.
We now define the vector field
\begin{equation}\label{solution.u}
	u(t,x) := \sum_{k=1}^\infty u_{k}(t,x) \,, \quad u_k(t,x):= \sum_{j=1}^{J_k} \overline{v_{k,j}}(t,x) + w_k(x)
\end{equation}
with pressure
\begin{equation}\label{pressure.u}
	p_{u}(t,x) :=  \sum_{k=1}^\infty p_{u_{k}}(t,x) \,, \quad p_{u_{k}}(t,x):= \sum_{j=1}^{J_k} \overline{p_{k,j}}(t,x) \,,
\end{equation}
where
\begin{equation}\label{vkj}
	\begin{aligned}
		\overline{v_{k,j}}(t,x) &:= \overline{v_{k}} (x'-\ty_{k,j}, x'' - \nu_{k,j}t)\,, \\ \overline{p_{k,j}}(t,x) & := \overline{p_{v_{k}}} (x'-\ty_{k,j}, x'' - \nu_{k,j}t)\,,
	\end{aligned}
\end{equation}
and
\begin{equation}\label{w.vf}
	w_k(x) = (0,F_k(x')) \,, \quad F_{k} : \T^m \to \R^{d-m}\,.
\end{equation}
Note that, no matter the choice of $F_k(x')$ sufficiently smooth is, the vector field $w_k(x)$ is a stationary solution with constant pressure of the Euler equations \eqref{euler.eq}, namely we have
\begin{equation}\label{wk.eq}
	w_k \cdot \nabla w_k = 0 \,, \quad {\rm div}\,w_k = 0 \,.
\end{equation}
To make sure that $u(t,x)$ is indeed a solution of \eqref{euler.eq}, we need to specify the functions $F_k(x')$. In particular, we choose
\begin{equation}\label{Fk}
	F_k(x') := \sum_{j=1}^{J_k} \nu_{k,j} \,\chi_{k} (|x'-\ty_{k,j}|) \,,
\end{equation}
where $\chi_{k}(r) \in \cC^{\infty}(\R)$ is an even cut-off function satisfying
\begin{equation}\label{cutoff.k}
\begin{aligned}
	& \chi_{k}(r) = 1 \quad \text{when} \ \ |r|< \varepsilon^k\,, \quad 	\chi_{k}(r) = 0 \quad \text{when} \ \ |r|> 2 \varepsilon^k\,, \quad \chi_{k} \in [0,1] \,, \\
	& |\partial_r^n \chi_k(r) | \leq C_n \varepsilon^{- k n}, \quad \forall \, r \in \R, \quad n \in \N \cup \{ 0 \}\,,
	\end{aligned}
\end{equation}
for some constant $C_n > 0$. For each $k\in\N$, we have that $F_k \in \cC^\infty(\T^m,\R^{d-m})$ and that the vector field $w_k(x)$ is locally equal to $(0,\nu_{k,j})$ when $x\in \spt(\overline{v_{k,j}})$. Note that each pair $(u_{k}(t,x),p_{u_k}(t,x))$ defined above by \eqref{solution.u}-\eqref{Fk} has actually the form of a quasi-periodic solution of the Euler equations \eqref{euler.eq}  as provided in \cite{EPsTdL}, which has been reproduced here on supports of scale $\varepsilon^{k}$. Moreover, by construction and by {\bf (A)}, the support in space of $(u_{k}(t,x),p_{u_k}(t,x))$ is in $\big(\bigcup_{j=1}^{J_k} B_{m,2\varepsilon^k}(\ty_{k,j}) \big)\times \T^{d-m}$ and it is disjoint from the one of $(u_{k'}(t,x),p_{u_{k'}}(t,x))$ for any $k'\neq k$. We use these properties to check that the pair $(u(t,x),p_{u}(t,x))$  in \eqref{solution.u}-\eqref{pressure.u} is indeed a solution of \eqref{euler.eq} as well.

First, we prove that each pair $(u_k(t,x), p_{u_k}(t,x))$ is a solution of \eqref{euler.eq} and we provide estimates on the seminorms.

\begin{lem}\label{lemma.uk}
	Assume that $\nu = (\nu_{k})_{k\in\N}$ satisfies \eqref{small.freq}.
	For each $k\in\N$, the vector field $u_k(t,x) $ is in $ \cC_{\rm div}^{\infty}(\T^d,\R^{d})$, with pressure $p_{u_{k}}(t,x)$ in $\cC^\infty(\T^d)$, is a solution of the Euler equations \eqref{euler.eq}, namely
	\begin{equation}\label{euler.uk}
		\pa_{t} u_k + u_k \cdot \nabla u_k + \nabla p_{u_k} = 0 \,, \quad {\rm div }\, u_k = 0\,,
	\end{equation}
	compactly supported in space in $\bigcup_{j=1}^{J_k} B_{m,2\varepsilon^k}(\ty_{k,j}) \times \T^{d-m}$. Moreover, we have the estimates, for any integer $n \geq 0$,
%	\begin{equation}\label{uk.esti}
%			\sup_{t\in\R} \|u_k(t,\cdot\,)\|_{n,\infty} \leq C_n \varepsilon^{k(S-n)-S} \,, \quad  	\sup_{t\in\R} \|p_{u_k}(t,\cdot\,)\|_{n,\infty} \leq C_n \varepsilon^{k(2S-n)-2S} \,,
%	\end{equation}
\begin{equation}\label{uk.esti}
	\begin{aligned}
		\sup_{t\in\R} \|u_k(t,\cdot\,)\|_{n,\infty} & \leq C_n \varepsilon^{k(S + 1-n)-S - 1} \,, \\
		\sup_{t\in\R} \|\pa_{t}u_k(t,\cdot\,)\|_{n,\infty} & \leq C_n \varepsilon^{k(2S + 2-(n+1))-2S - 2} \,, \\
		 \sup_{t\in\R} \|p_{u_k}(t,\cdot\,)\|_{n,\infty}& \leq C_n \varepsilon^{k(2S+ 2-n)-2S - 2} \,,
	\end{aligned}
\end{equation}
	for some constant $C_n>0$ independent of $\varepsilon\in (0,1)$ and of $k\in\N$.
\end{lem}
\begin{proof}
	By \eqref{solution.u}-\eqref{Fk}, Proposition \ref{gravi.sol} and by {\bf (A)}, we compute
	\begin{equation}\label{ditu}
		\begin{aligned}
				\pa_{t} u_k & = - \sum_{j=1}^{J_k} \nu_{k,j} \cdot \nabla'' \overline{v_{k}}(x'-\ty_{k,j},x''-\nu_{k,j} t) \,,
		\end{aligned}
	\end{equation}
	and, using \eqref{ditu},
	\begin{equation}
		\begin{aligned}
			u_k \cdot \nabla u_k & = \Big( \sum_{j=1}^{J_k} \overline{v_{k,j}} + w_k \Big) \cdot  \Big( \sum_{j=1}^{J_k}\nabla \overline{v_{k,j}} + \nabla w_k \Big) \\
			& = \sum_{j=1}^{J_k} \overline{v_{k,j}}\cdot \nabla \overline{v_{k,j}} + \sum_{j=1}^{J_{k}} \big(  \overline{v_{k,j}} \cdot \nabla w_k + w_k \cdot \nabla \overline{v_{k,j}} \big) + w_k \cdot \nabla w_k \\
			& = -  \sum_{j=1}^{J_k}  \nabla \overline{p_{k,j}} + \sum_{j=1}^{J_k} \big(0+ \nu_{k,j}\cdot \nabla'' \overline{v_{k}} (x'-\ty_{k,j},x''-\wt\nu_{k,j}t) \big)  +0 \\
			& = - \nabla p_{u_k} - \pa_{t} u_k \,.
		\end{aligned}
	\end{equation}
	This, together with the fact that ${\rm div}\,\overline{v_{k,j}} =0$ for any $j=1,...,J_k$ by \eqref{vkj} and Proposition \ref{gravi.sol}, concludes the proof of \eqref{euler.uk}. \\
	To prove the estimates \eqref{uk.esti}, we first note that, by \eqref{cutoff.k}, we have $\|\chi_{k}\|_{n,\infty} \leq C_n \varepsilon^{-nk}$ for any integer $n \geq 0$. Moreover, by \eqref{small.freq}, we have that $|\nu_{k}|\leq C \varepsilon^{(S + 1)(k-1)}$, for some constant $C>0$ independent of $k\in\N$. Therefore, we deduce, that $w_k(x)$ in \eqref{w.vf}-\eqref{Fk} satisfies $\|w_{k}\|_{n,\infty} \leq C_n \varepsilon^{(S + 1-k)n-S - 1}  $  for any integer $n\geq 0$, recalling that $(J_k)_{k\in\N}\in \ell^\infty(\N,\N)$. Furthermore, by \eqref{ditu} and using the fact that each $\overline{v_{k,j}}(t,x)$ is supported in space on the cylinder $B_{m,2\varepsilon^k}(\ty_{k,j})\times \T^{d-m}$, disjoint for any $j'\neq j$  from the cylinder $ B_{m,2\varepsilon^k}(\ty_{k,j'}) \times \T^{d-m}$ supporting $\overline{v_{k,j'}}(t,x)$, we estimate, for any integer $n \geq 0$ and uniformly in $t\in\R$,
	\begin{equation}\label{stima.ditu}
	\begin{aligned}
			\| \pa_{t}u_{k}(t,\,\cdot) \|_{n+1, \infty} & \sup_{j=1,...,J_k}|\nu_{k,j}| \| \nabla'' \overline{v_{k}}(x'-\ty_{k,j},x''-\wt\nu_{k,j} t)   \|_{n,\infty} \\
			 &\lesssim  \varepsilon^{(S+1)(k-1)} \| v_{k} \|_{n+1,\infty}  \\
			 & \lesssim_{n}   \varepsilon^{(S + 1)(k-1)} \varepsilon^{k(S + 1-(n+1))-S - 1} \,.
	\end{aligned}
	\end{equation}
	 Collecting together \eqref{solution.u}, \eqref{pressure.u}, \eqref{w.vf} and estimates \eqref{rescaled.esti}, \eqref{stima.ditu}, we obtain the estimates \eqref{uk.esti} and the proof is concluded.
\end{proof}

We now show that $u(t,x)$ in \eqref{solution.u} solves the Euler system \eqref{euler.eq} and that it has the desired regularity and estimates. 

\begin{prop}
	The vector field $u(t,x)$ in \eqref{solution.u}, with pressure $p_{u}(t,x)$ as in \eqref{pressure.u}, is a solution of the Euler equations \eqref{euler.eq}. Moreover, assuming that $\nu = (\nu_{k})_{k\in\N}$ satisfies \eqref{small.freq}, we have $u(t,\,\cdot\,) \in \cC_{\rm div}^{S + 1}(\T^d,\R^d)$, $\partial_t u(t, \cdot) \in \cC_{\rm div}^{2 S + 1}(\T^d,\R^d)$ and $p_{u}(t, \cdot) \in \cC^{2 S + 2}(\T^d)$, with estimates, 
	\begin{equation}\label{u.final.esti}
		\begin{aligned}
		\sup_{t\in\R} \|u(t,\cdot\,)\|_{n,\infty}& \leq C_n \varepsilon^{-S - 1} \,, \quad \forall\, n=0,1,...,S + 1\,, \\
		\sup_{t\in\R} \|\pa_{t} u(t,\cdot\,)\|_{n,\infty}& \leq C_n \varepsilon^{-2S - 2} \,, \quad \forall\, n=0,1,...,2S + 1\,, \\
		\sup_{t\in\R} \|p_{u}(t,\cdot\,)\|_{n,\infty}& \leq C_n \varepsilon^{-2S - 2}\,, \quad \forall\, n=0,1,...,2S + 2 \,.
		\end{aligned}
	\end{equation}
\end{prop}
\begin{proof}
	By Lemma \ref{lemma.uk} and by {\bf (A)}, each vector field of the sequence $(u_k(t,x))_{k\in\N}$ is compactly supported in space and all these supports are pairwise disjoint. We use this properties and the fact that each $u_k(t,x)$ in solves \eqref{euler.uk} to compute, with $u(t,x)$ and $p_u(t,x) $ as in  \eqref{solution.u}, \eqref{pressure.u},
	\begin{equation}
		\begin{aligned}
			 u\cdot \nabla u &= \sum_{k=1}^\infty u_k \cdot \sum_{k=1}^\infty \nabla u_k = \sum_{k=1}^\infty u_k \cdot \nabla u_k = \sum_{k=1}^\infty \big( - \nabla p_{u_k}  - \pa_{t}u_k\big) = - \nabla p_{u} -\pa_{t} u \,, \\
			{\rm div} \, u & =  \sum_{k=1}^\infty {\rm div} \, u_k = 0 \,,
		\end{aligned}
	\end{equation}
	which indeed proves \eqref{euler.eq}. It remains to prove the finite regularity of the solution. By \eqref{solution.u}, \eqref{uk.esti}, since the support in space are pairwise disjoint, we have that, for any $n=0,1,...,S + 1$,
	\begin{equation}
		\sup_{t\in\R} \|u(t,\cdot\,)\|_{n,\infty} = \sup_{t\in\R} \sup_{k\in\N} \|u_k(t,\,\cdot\,)\|_{n,\infty} \leq C_n \varepsilon^{-S - 1} \sup_{k\in\N} \varepsilon^{k(S + 1-n)} \leq C_n \varepsilon^{-S - 1}\,.
	\end{equation}
	The estimates for $ \partial_t u(t, \cdot), p_u(t, \cdot)$ can be proved similarly and we omit them. Hence \eqref{u.final.esti} follows. This concludes the proof.
\end{proof}

In order to conclude the proof of Theorem \ref{teorema}, it remains to show the existence of the embedding $U:(\T^{N_k})_{k\in\N} \to \cC_{\rm div}^S(\T^d,\R^d)$. 
%For any $k\in\N$, recalling that $N_k\leq (d-m)J_k$, we consider a linear embedding $\cN_{k}:\T^{N_k} \to \T^{(d-m)J_k}$ of the form
%\begin{equation}
%	\T^{N_k}\ni \vartheta_{k} \to \cN_{k}(\vartheta_{k}):=\big( \cN_{k,1}(\vartheta_{k}),..., \cN_{k,J_k} (\vartheta_{k})\big) \,,\quad \cN_{k,j}: \T^{N_k} \to \T^{d-m}\,,
%\end{equation}
%where the linearity is meant at the level of the lift to the universal cover $\R^{(d-m)J_{k}}$. Let $ \overline{\cN}_{k}:= \di_{\vartheta_{k}} \cN_{k}(0):\R^{N_k} \to \R^{(d-m)J_k}$. We now define the map $\cN : (\T^{N_k})_{k\in\N} \to (\T^{(d-m)J_k})_{k\in\N}$ by 
%\begin{equation}
%	(\T^{N_k})_{k\in\N} \ni (\vartheta_{k})_{k\in \N} = \vartheta \mapsto \cN(\vartheta) = \big(\cN_{k}(\vartheta_{k}) \big)_{k\in\N} \in (\T^{(d-m)J_k})_{k\in\N} \,.
%\end{equation}
%Abusing notation, we denote 
%%\begin{equation}
%%	\overline{\cN} := (\overline{\cN}_k)_{k\in\N} := \di_{\theta} \cN(0) : (\R^{N_k})_{k\in\N} \to (\R^{(d-m)J_k})_{k\in\N}\,,
%%\end{equation}
%\begin{equation}
%	 \di_{\vartheta} \cN(0)=:(\overline{\cN}_k)_{k\in\N}  : (\R^{N_k})_{k\in\N} \to (\R^{(d-m)J_k})_{k\in\N}\,,
%\end{equation}
%where the differential $\di_{\vartheta}$ is meant in the sense of Fr\'echet. 
We  define the claimed family of (almost-periodic) solutions as $\theta\to U(\theta+ \nu t)$, where  the embedding $U: (\T^{N_k})_{k\in\N}\to \cC_{\rm div}^S(\T^d,\R^d)$ is given by
\begin{equation}
	\begin{aligned}
		u(t,x) &:= U(\theta+\nu t)(x) = \sum_{k\in\N} u_k(t,x) = \sum_{k\in \N} U_k(\theta_{k} + \nu_{k} t)(x) \,, 
	\end{aligned}
\end{equation}
with each $ U_k: \T^{N_k}\to \cC_{\rm div}^\infty(\T^d,\R^d) \subset \cC_{\rm div}^S(\T^d,\R^d)$. for $k\in\N$, given by
\begin{equation}
	\begin{aligned}
		& U_k(\theta_{k} + \nu_{k} t)(x)  := \sum_{j=1}^{J_k} \overline{v_k} \big( x' - \ty_{k,j} , x'' - \theta_{k,j}- \nu_{k,j} t \big) + \sum_{j=1}^{J_k} \big( 0, \nu_{k,j}\, \chi_{k}(|x'-\ty_{k,j}|) \big)\,, \\
		& \theta_k = (\theta_{k, j})_{j = 1, \ldots, J_k} \in \T^{N_n}, \quad \theta_{k, j} \in \T^{d - m}, \quad \forall \,k \in \N\,,\ \  j = 1, \ldots, J_k \,,
	\end{aligned}
\end{equation}
(recall that $N_k = (d - m) J_k$, see \eqref{sequence.Nk}) and  with initial data
%\begin{equation}\label{initial.data}
	\begin{align}
		u_{\theta}(x) &:= U(\theta)(x) = \sum_{k\in\N} u_{\theta_{k}}(x) = \sum_{k\in \N} U_k(\theta_{k})(x) \,,  \label{initial.data} \\
		U_k(\theta_{k})(x) & := \sum_{j=1}^{J_k} \overline{v_k} \big( x' - \ty_{k,j} , x'' - \theta_{k,j}  \big) + \sum_{j=1}^{J_k} \big( 0,  \nu_{k,j}\, \chi_{k}(|x'-\ty_{k,j}|) \big)\,.
	\end{align}
%\end{equation}

As last step, we prove the estimate on the continuity and the differentiability of the embedding $U$.
\begin{prop}
	Assume that $\nu = (\nu_{k})_{k\in\N}$ satisfies \eqref{small.freq}.
	Then the embedding $U: (\T^{N_k})_{k\in\N}\to \cC_{\rm div}^S(\T^d,\R^d)$ is $\cC_b^1$ according to Definition \ref{C1-smooth.def}, with estimates
%		\begin{equation}\label{last.esti}
		\begin{align}
			& \sup_{\vartheta \in (\T^{N_k})_{k\in\N}} \|  U(\vartheta) \|_{n,\infty} \leq C_n  \varepsilon^{-S-1} \,, \quad   0\leq n\leq S\,, \label{last.esti}  \\
			&	\sup_{\vartheta \in (\T^{N_k})_{k\in\N}} \| \di_{\vartheta} U(\vartheta) [\wh\vartheta] \|_{n,\infty} \leq C_n \varepsilon^{-S-1} | \wh\vartheta |_{\infty}  \quad \forall\,\wh \vartheta \in (\R^{N_k})_{k\in\N}\,,  \quad 0\leq n\leq S \,.
		\end{align}
%	\end{equation}
\end{prop}
\begin{proof}
	The first estimate in \eqref{last.esti} follows by \eqref{initial.data} and \eqref{u.final.esti}. We now prove the second estimate in \eqref{last.esti}. By \eqref{initial.data}, we compute, for any $\vartheta\in (\T^{N_k})_{k\in\N}$ and $\wh\vartheta \in (\R^{N_k})_{k\in\N}$,
	\begin{equation}\label{natale1}
	\di_{\vartheta} U(\vartheta)[\wh\vartheta] = - \sum_{k\in\N} \sum_{j=1}^{J_k} \wh\vartheta_{k} \cdot \nabla'' \overline{v_k} (x'-\ty_{k,j}, x'' - \vartheta_{k,j}) \,.
	\end{equation}
	Therefore, by \eqref{rescaled.esti}, using the fact that each term in the series in \eqref{natale1}  is supported in space on the cylinder $B_{m,2\varepsilon^k}(\ty_{k,j})\times \T^{d-m}$, that these supports are  disjoint one from the other, and that $|\wh\vartheta_{k}|\leq |\wh\vartheta|_{\infty}$ , we obtain, for any $\vartheta\in(\T^{N_k})_{k\in\N}$, $\wh\vartheta \in (\R^{N_k})_{k\in\N}$ and for any $n=0,1,...,S$,
%	\begin{equation}
%	\begin{aligned}
%			\| 	\di_{\vartheta} U(\vartheta)[\wh\nu] \|_{n,\infty} & \leq \sum_{k\in\N} \sum_{j=1}^{J_k} \| \nabla'' \overline{v_k} \|_{n,\infty} |\wh\nu|_\infty   \leq \| (J_k) \|_{\ell^\infty}  \sum_{k\in \N} \| \overline{v_k} \|_{n+1,\infty} |\wh\nu|_{\infty} \\
%			& \leq C_n \sum_{k\in \N} \varepsilon^{k(S + 1-(n+1))-S  - 1} |\wh\nu|_\infty \leq C_n \varepsilon^{-S} |\wh\nu|_\infty\,.
%	\end{aligned}
%	\end{equation}
	\begin{equation}
	\begin{aligned}
			\| 	\di_{\vartheta} U(\vartheta)[\wh\vartheta] \|_{n,\infty} & \leq \sup_{k\in\N} \sup_{j=1,...,J_k} \| \nabla'' \overline{v_k}(\,\cdot\, - \ty_{k,j}, \,\cdot\, -\vartheta_{k,j}) \|_{n,\infty} |\wh\vartheta|_\infty  \\ 
		 & 	\leq \sup_{k\in\N} \sup_{j=1,...,J_k} \|  \overline{v_k}(\,\cdot\, - \ty_{k,j}, \,\cdot\, -\vartheta_{k,j}) \|_{n+1,\infty} |\wh\vartheta|_\infty  \\ 
		& \leq C_n \sup_{k\in\N} \varepsilon^{k(S + 1-(n+1))-S  - 1} |\wh\vartheta|_\infty \leq C_n \varepsilon^{-S-1} |\wh\vartheta|_\infty\,.
	\end{aligned}
\end{equation}
	This implies the claimed estimate and concludes the proof.
\end{proof}

\begin{footnotesize}
	
\end{footnotesize}

\bigskip

\begin{flushright}
	\textbf{Luca Franzoi}
	
	\smallskip
	
	Dipartimento di Matematica ``Federigo Enriques''
	
	Universit\`a degli Studi di Milano
	
	Via Cesare Saldini 50
	
	20133 Milano, Italy
	
	\smallskip 
	
	\texttt{luca.franzoi@unimi.it}
	
\end{flushright}
\begin{flushright}
	\textbf{Riccardo Montalto}
	
	\smallskip
	
	Dipartimento di Matematica ``Federigo Enriques''
	
	Universit\`a degli Studi di Milano
	
	Via Cesare Saldini 50
	
	20133 Milano, Italy
	
	\smallskip 
	
	\texttt{riccardo.montalto@unimi.it}
	
\end{flushright}

 \end{document}